Find b^2-4ac adn the number of real solutions to this equation:
9m^2 + 16 = 24m
I am confused !!! Can you help me? :-)
Thanks,

Answer by RAY100(1637)

Let D = Discriminant = (b^2 -4 ac)
.
basic form is ax^2 +bx+c =0
.
when D=0,,,one solution, line touches x axis one time
.
when D>0,,,two solutions, line intercepts x axis 2 times
.
when D<0,,,no (REAL) solutions, line does not touch x axis
.
in this case,, 9m^2 +16 = 24m
.
9m^2 -24m +16 =0
.
D= { (-24)^2 -4(9)(16)} = {576 - 576} =0,,,,,and one solution
.
solving for m using the quadratic equation demonstrates the use of D
.
m=[-b +/- sqrt D] /2(a) = {-(-24) +/- sqrt(0) }/2(9)= 24/18 = 4/3=1 1/3
.
NOTE: @ D=0, m= {(-b) +/- 0 }/2a = -b/2a,,,,or one number
.
@ D >0,,m= { (-b) +/- sqrt(+#) } /2a,,,or two #'s
.
@ D <0,,,m= { (-b) +/- sqrt(-#) }/2a = complex number
.
to check : factor,,, 9m^2 -24m +16,,(3m-4)(3m-4),,
.
set equal to zero,,,,3m-4=0,,,m=4/3 (twice),,,,but only one solution
.